# Promptus: Can Prompts Streaming Replace Video Streaming with Stable Diffusion

### v1

@@ 4.1 The sender side
Initialization for the first frame. As stated in §3.1, to improve the quality of the generated images, Promptus uses the noisy previous frame as the input noise for Stable Diffusion. However, for the first frame of the video (the frame index starts from 1), there is no previous frame $Z^{0}$. To address this, Promptus uses the VAE’s Encoder to map the first frame itself to the latent space, obtaining $Z^{0}$. Then, noise is added to $Z^{0}$ according to Equation 1. The noisy $N^{1}$ will be used as the input noise for Prompt Inversion of the first frame. Since the purpose of using the noisy previous frame is to reduce the distance between the input noise and the target image in the latent space, the noisy current frame naturally works as well. Note that this noise $Z^{0}$ will also be sent to the receiver side along with the prompt for generating the first frame image. On the other hand, due to the absence of a previous frame, Promptus does not apply temporal smoothing regularization for the Fitting of the first frame.
@@
Sparse prompt streaming. As stated in §3.3, through temporal smoothing regularization, the prompts of most frames can be approximated by linear interpolation of the prompts of keyframes. Therefore, Promptus defines the keyframe interval as $K$ (the choice of $K$ is discussed in §5.2), adding one keyframe every $K$ frames. Only the prompts of keyframes will be streamed to the receiver. Note that when the aforementioned abrupt scene changes occur, the difference between frames in the latent space is significant, and the interpolation of prompts no longer works. In this case, the last frame before the scene change will also become a keyframe and be sent. The frames after the scene change will start a new count.
Low-rank based adaptive bitrate. According to §3.2, the higher the bitrate of the prompt, the higher the quality of the generated image. Therefore, in prompt streaming, it is necessary to increase the prompt bitrate as much as possible while avoiding network congestion to maximize the user experience. Promptus adopts the WebRTC [1] framework for prompt streaming, which dynamically probes the currently available bandwidth. Consequently, Promptus adaptively adjusts the rank to make the prompt bitrate close to the probed bandwidth. It is worth noting that the prompts transmitted by Promptus are not encoded. Therefore, compared to codec-based streaming, Promptus can precisely control the bitrate.
@@ 4.2 The receiver side
Video generation based on sparse prompts. After receiving the prompt of keyframe $i$, the receiver first performs linear interpolation between the prompts of adjacent keyframes $i-k$ and $i$ to approximate the prompts of the intermediate $K-1$ frames. Afterward, these prompts are used for Stable Diffusion’s image generation. As described in §3.1, the generated $Z$ for each frame will be noised and used as the input noise for the generation of the next frame. Note that since the ratio of the noise (95%) in the noisy previous frame is much larger than the ratio of the previous frame (5%), the error or complete loss of the previous frame (earlier frames can be used) has almost no impact on the generation of subsequent frames.
Real-time video generation. The frame rate of video playback at the receiver depends on Promptus’s image generation speed, making real-time generation crucial. Promptus’s image generation consists of five parts: prompt dequantization, prompt composition, prompt interpolation, adding noise to previous frame and Stable Diffusion image generation. The first four parts only involve simple linear calculations, so their time consumption can be ignored. The speed of Stable Diffusion becomes the bottleneck. To address this, we adopt the StreamDiffusion [28], which accelerates Stable Diffusion image generation to 100 FPS through batch processing. By this, Promptus achieves real-time video generation on the receiver.
@@ 5 Evaluation
Our four main evaluation results are as follows: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases. Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines. Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
@@ 5.1 Experiment Setup
Test videos: To validate the generalizability of Promptus across different video domains, we selected 7 videos from 4 datasets with vastly different content, as summarized in Table 1. Specifically, the domains of these videos span natural landscapes and human activities, outdoor long-range scenes and indoor close-up scenes, real-world scenes and CG-synthesized scenes, 3D gaming scenes and 2D animations. All videos are cropped and resized to a resolution of 512*512, with a frame rate of 30 FPS. Since Promptus uses a pre-trained and frozen Stable Diffusion, it does not involve model training and does not require training videos. However, for fairness, we still checked that the above test videos were not used during the pre-training of Stable Diffusion.
Table 1: Test videos summary
Dataset #Videos #frames Description
QST [53] 2 300 Natural landscapes, outdoor distant view
UVG [32] 2 300 Human activity, face, hand, indoor close-up
GTA-IM [17] 2 300 3D Game recording, CG-synthesized scenes
Animerun [43] 1 60 2D animation, cartoon
Total 7 960
Baselines: We compare Promptus with two baselines: H.265 [5] and VAE [27, 38]. H.265 is an advanced traditional codec that achieves compression by utilizing hand-designed intra-frame prediction, motion compensation, and transform coding techniques. Moreover, H.265 can control the bitrate of the encoded video by adjusting the quantization parameters, balancing quality and bitrate. Since it does not involve learnable parameters, it has good generality. VAE is a deep learning-based neural codec. Both its encoder and decoder are trainable neural networks. The encoder maps the input image to a low-dimensional latent variable, while the decoder reconstructs the image from the latent variable. The dimensionality of the latent variable is usually much smaller than that of the original image, thus achieving compression. For inter-frame compression of VAE, we first quantize the latent variables of each frame and then encode them into a video using H.265. The quantization process causes almost no quality degradation, and the quality degradation is mainly due to video encoding. Therefore, we adjust the target bitrate of video encoding to balance the quality and bitrate of VAE. For fairness, the training set of VAE is the same as that of Stable Diffusion, both consisting of 5.85 billion images [6].
Metric: To evaluate video quality, we adopt the LPIPS [55] (Learned Perceptual Image Patch Similarity) instead of the traditional SSIM and PSNR. This is because LPIPS has a significantly higher correlation with human subjective ratings compared to SSIM and PSNR [51, 55], which means that LPIPS better reflects human subjective perception of video quality. LPIPS uses a pre-trained neural network to extract layer-wise features from the original image and the distorted image, and then calculates the normalized L2 distance between the features as the LPIPS value. Therefore, a smaller LPIPS value indicates a higher quality of the distorted image. Specifically, we use VGG [41] pre-trained on ImageNet [18] as the feature extractor for LPIPS.
@@
Network trace: We collected 6 network traces from real-world scenarios such as subways, driving, and walking, under 2G, 3G, and 4G networks. The traces consisted of 1 from a 4G network, 2 from 3G, and 3 from 2G. Each trace lasted between 5 and 30 seconds, totaling over 100 seconds. Our traces aimed to test various weak network conditions such as poor signal coverage, network overload, high-speed movement, and frequent switching of mobile communication networks. The average bandwidth per second ranged between 50 kbps and 4000 kbps, while the one-way network delay was approximately 30 ms to 100 ms.
Testbed: To test the streaming performance, we use WebRTC [1] to deploy Promptus and the baselines. WebRTC is a real-time video communication framework that dynamically probes the currently available bandwidth to avoid congestion. To simulate real network conditions, we use Mahimahi [34] to replay the above network traces. The queue length of Mahimahi is set to 60, and the drop-tail strategy is adopted.
@@ 5.2 Trade off between bitrate and quality
Figure 7: Visualization of the fitting results at different ranks. It can be seen that as the rank increases, the prompt can fit more details. When the rank is 4, the earrings are lost. When the rank increases to 16, the earrings are successfully fitted.
Prompt rank. We illustrate the variation in visual quality of Promptus under different ranks, as depicted in Figure 8. It indicates that the higher the prompt rank, the higher the quality of Promptus. For example, when the prompt rank increases from 4 to 16, the LPIPS decreases from 0.265 to 0.221 (on the line with a keyframe interval of 1). This is because the larger the rank, the stronger the representational capability of the prompt, allowing it to fit the target image more accurately, as described in §3.2. We present a visualization of the fitting results at different ranks, as shown in Figure 7. It can be seen that as the rank increases, the prompt can fit more details. For instance, when the rank is 4, the earrings are lost in the fitting result, while when the rank increases to 16, the earrings are successfully fitted.
Keyframe interval. Figure 8 also shows the impact of different keyframe intervals on quality. The smaller the keyframe interval, the higher the quality of Promptus. For example, when the prompt rank is 16, reducing the keyframe interval from 10 to 1 decreases the LPIPS from 0.274 to 0.221. This is because, when the interval is smaller, the distance between keyframes in the prompt space is smaller. Since we constrain the prompt space to be temporally smooth, the linear interpolation of keyframe prompts can more accurately approximate the prompts of intermediate frames, as described in §3.3.
Figure 8: The impact of prompt rank and keyframe interval on visual quality. It indicates that visual quality improves with increasing prompt rank and decreasing keyframe interval.
Quality-bitrate tradeoff. Increasing the prompt rank and reducing the keyframe interval lead to a rapid rise in bitrate while improving quality. Therefore, we present the tradeoff between bitrate and quality for Promptus, as shown in Figure 9. First, it illustrates that overall, the quality monotonically increases with the increase in bitrate. Thus, to optimize quality, Promptus sends prompts at a bitrate closest to the available bandwidth. Second, Promptus can achieve scalable bitrates. For example, by adjusting the rank in the range of 4 to 32 and the keyframe interval in the range of 2 to 8, Promptus’s bitrate spans from 113 kbps to 4284 kbps. Third, at the same bitrate, the quality varies for different configurations. For instance, when the bitrate is 550 kbps, the configuration with a rank of 8 and a keyframe interval of 4 has an LPIPS of 0.255, while the configuration with a rank of 16 and a keyframe interval of 8 has an LPIPS of 0.264, which is lower in quality than the former. So at the same bitrate, Promptus tends to choose configurations with smaller keyframe intervals for streaming.
Figure 9: The tradeoff between bitrate and quality for Promptus. The quality monotonically increases with the increase in bitrate. Besides, Promptus can achieve scalable bitrates.
5.3 Compression efficiency
This section demonstrates the compression efficiency of Promptus. Figure 10 shows the CDF of the frame quality for Promptus and baselines under 4 bitrate levels. Since a lower LPIPS represents better visual quality, a leftward shift of the curve in the figure represents more high-quality frames, and thus higher compression efficiency. We also calculate the average LPIPS for each method, represented by the vertical lines in the figure.
@@ 5.3 Compression efficiency
Figure 10: Frame quality CDF and mean of Promptus and baselines at four target bitrates. It indicates Promptus achieves better compression efficiency across all bitrate levels. Besides, the lower the bitrate, the greater the advantage of Promptus.
@@
First, Promptus achieves better compression efficiency across all bitrate levels. For example, in Figure 10, the curves of Promptus are all to the left of the baseline curves. Second, the lower the bitrate, the greater the advantage of Promptus. At a high level, as the bitrate decreases from 540 kbps to 140 kbps, the distance between Promptus’s curves and the baselines’ curves gradually widens. At a low level, when the bitrate is 540 kbps, the average LPIPS of Promptus is 0.018 and 0.085 lower than VAE and H.265, respectively. When the bitrate is 140 kbps, this difference further increases to 0.139 and 0.118. This is because when the bitrate is reduced, H.265 uses coarser quantization and loses many high-frequency details, resulting in blurriness and block artifacts in the video, which significantly impairs the perceptual quality. VAE mitigates this phenomenon by mapping images to a low-dimensional latent space, reserving more bitrate for video coding. Therefore, at most bitrates, the quality of VAE is superior to H.265. However, at extremely low bitrates (such as 140 kbps), the latent space inevitably introduces distortions caused by video coding. At this point, the VAE Decoder introduces a large number of errors when reconstructing the images, causing a significant decrease in VAE’s quality. On the other hand, when the bitrate is reduced, Promptus reduces the representational capability of the prompt rather than degrading the video quality. This prevents Promptus from accurately describing the video content, resulting in slight misalignments in the generated frames. However, thanks to the inherent image generation capability of Stable Diffusion, Promptus’s frames still have good sharpness and details, and thus the perceptual quality is better than VAE and H.265.
5.4 Generality
@@
This section demonstrates the generality of Promptus across different domains. Figure 12 shows the mean of the frame quality for Promptus and two baselines on four different datasets (as described in Table 1).
Figure 11: Visualization of the compression results on different datasets. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
Figure 12: Mean frame quality of Promptus and baselines on 4 different datasets. It demonstrates the generality of Promptus across different domains. Besides, the more high-frequency details a video has, the greater the advantage of Promptus.
@@
First, Promptus achieves better compression efficiency on each dataset. This is because, as shown in Figure 12, Promptus achieves lower average LPIPS compared to the baselines on each dataset. To intuitively demonstrate this improvement, we also visualize the compression results of the three methods on the four datasets at lower bitrates (such as 225 kbps), as shown in Figure 11. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
@@
Second, the more high-frequency details a video has, the greater the advantage of Promptus. For example, for the Animerun dataset with fewer details, the LPIPS of Promptus is 0.015 lower than H.265, which is not a significant advantage. However, for the detail-rich UVG, this difference further expands to 0.121. This is because for 2D animations with large areas of solid colors and simple details, H.265’s intra-frame prediction, block partitioning, and motion compensation techniques can handle them well, so Promptus’s performance gain is small. For detail-rich real-world videos, H.265 discards more high-frequency information during compression, thus damaging the perceptual quality (especially at lower bitrates). Although Promptus also loses high-frequency information from the original image, Stable Diffusion completes the lost high-frequency information based on prior knowledge during generation, resulting in a smaller perceptual quality loss.

### v2

@@ 4.1 The sender side
Initialization for the first frame. As stated in §3.1, to improve the quality of the generated images, Promptus uses the noisy previous frame as the input noise for Stable Diffusion. However, for the first frame of the video (the frame index starts from 1), there is no previous frame $Z^{0}$. To address this, Promptus uses the VAE’s Encoder to map the first frame itself to the latent space, obtaining $Z^{0}$. Then, noise is added to $Z^{0}$ according to Equation 1. The noisy $N^{1}$ will be used as the input noise for Prompt Inversion of the first frame. Since the purpose of using the noisy previous frame is to reduce the distance between the input noise and the target image in the latent space, the noisy current frame naturally works as well. Note that this noise $Z^{0}$ will also be sent to the receiver side along with the prompt for generating the first frame image.
@@
Sparse prompt streaming. As stated in §3.3, through interplation-aware fitting, the prompts of most frames can be approximated by linear interpolation of the prompts of keyframes. Therefore, Promptus defines the keyframe interval as $K$ (the choice of $K$ is discussed in §5.2), adding one keyframe every $K$ frames. Only the prompts of keyframes will be streamed to the receiver. Note that when the aforementioned abrupt scene changes occur, the difference between frames in the prompt space is significant, and the interpolation of prompts no longer works. In this case, the last frame before the scene change will also become a keyframe. The frames after the scene change will start a new count.
Low-rank based adaptive bitrate. According to §3.2, the higher the bitrate of the prompt, the higher the quality of the generated image. Therefore, in prompt streaming, it is necessary to increase the prompt bitrate as much as possible while avoiding network congestion to maximize the user experience. Consequently, Promptus fits multiple prompts of different ranks in advance. During streaming, Promptus selects the prompt with the bitrate closest to the estimated bandwidth for transmission. It is worth noting that the prompts transmitted are not entropy-encoded. Thus, compared to codec-based streaming, Promptus can precisely control the bitrate.
@@ 4.2 The receiver side
Video generation based on sparse prompts. As shown in Figure 6, after receiving the prompt of keyframe $i$, the receiver first performs linear interpolation between the prompts of adjacent keyframes $i-k$ and $i$ to approximate the prompts of the intermediate $K-1$ frames. Afterward, these prompts are used for Stable Diffusion’s image generation. As described in §3.1, the generated $Z$ for each frame will be noised and used as the input noise for the next frame’s generation. Note that since the ratio of the noise (95%) in the noisy previous frame is much larger than the ratio of the previous frame (5%), the error of the previous frame has almost no impact on the generation of subsequent frames.
Real-time video generation. The frame rate of video playback at the receiver depends on Promptus’s image generation speed, making real-time generation crucial. Specifically, Promptus’s generation consists of five parts: prompt dequantization, prompt composition, prompt interpolation, adding noise to previous frame and Stable Diffusion generation. The first four parts only involve simple linear calculations, so their time consumption can be ignored. The speed of Stable Diffusion becomes the bottleneck. To address this, following StreamDiffusion (kodaira2023streamdiffusion, ), we first replace the VAE in Stale Diffusion with a more lightweight TAESD. To further accelerate generation, instead of using PyTorch, we adopt TensorRT to build inference engines for Promptus. In this way, Promptus achieves real-time video generation at the receiver. A detailed evaluation of generation speed can be found in §5.9.
@@ 5.1 Experiment Setup
Test videos: To validate the generality across different domains, we test Promptus on four video datasets with vastly different content: QST zhang2020dtvnet , UVG mercat2020uvg , GTA-IM caoHMP2020 and Animerun siyao2022animerun . The domains of these videos span natural landscapes and human activities, outdoor long-range scenes and indoor close-up scenes, real-world scenes and CG-synthesized scenes, 3D video games and 2D animations. For consistency, videos are cropped and resized to a resolution of 512*512, with a frame rate of 30 FPS. Note that Promptus supports flexible resolutions and we test other resolutions in §5.6.
Baselines: We compare Promptus with four baselines: H.265 (265, ), VAE (rombach2022high, ; kingma2014auto, ), NAS (yeo2018neural, ) and VQGAN (esser2021taming, ; li2023reparo, ):
  • H.265 is an advanced traditional codecs that achieve compression by utilizing hand-designed modules.
  • VAE is a learning-based neural codec. Both its encoder and decoder are trainable neural networks. The encoder maps the input image to low-dimensional latent variables, while the decoder reconstructs the image from the latent variables. The dimensionality of the latent variables is usually much smaller than that of the original image, thus achieving compression. For inter-frame compression, we encode the latent variables using H.265.
  • NAS represents neural-enhanced streaming, which trains overfitted super-resolution DNNs for each video. These DNNs are transmitted along with low-resolution videos. At the receiving end, DNNs are employed to enhance the quality of low-resolution videos. To control the total bitrate, we adjust the bitrate of the low-resolution video and the number of DNN parameters transmitted.
  • VQGAN stands for generative streaming. To compress the video, it first divides the video into patches, maps each patch to a latent variable using a VAE, and then quantizes each latent variable into a 10-bit token index using a codebook. To support general videos, the codebook is trained on diverse visual domains plummer2015flickr30k , not limited to facial videos (like Reparo (li2023reparo, )). At the receiver, the codebook is shared, so the token indices can be reconstructed into the video. To control the bitrate, we adjust the number of patches (token indices) for streaming.
Figure 7: Visualization of the fitting results at different ranks. It can be seen that as the rank increases, the prompt can fit more details. For example, when the rank is 4, the earrings in (a) are lost, and the lamp’s color in (d) is inconsistent with the ground truth. When the rank increases to 16, the earrings in (b) are successfully fitted, and the lamp’s color in (e) is also corrected.
Metric: To evaluate video quality, we mainly adopt the LPIPS (zhang2018unreasonable, ), because Promptus is more oriented towards the semantic communication task rather than video compression. Our primary goal is to keep the semantic information correct at low bitrates. LPIPS better reflects human subjective perception quality (yu2023language, ; zhang2018unreasonable, ), making it more suitable for evaluating semantic distortion. Smaller LPIPS indicates higher quality. We also evaluate traditional PSNR and SSIM as objective metrics, with discussions provided in §6.
@@
Network trace: We collect network traces from real-world scenarios such as subways, driving, and walking, under 2G, 3G, and 4G networks. Our traces aim to test various weak network conditions such as poor signal coverage, network overload, high-speed movement, and frequent switching of mobile wireless networks. The average bandwidth per second ranges between 50 kbps and 4000 kbps, while the one-way network delay is about 30 ms to 100 ms. We use Mahimahi (netravali2015mahimahi, ) to replay the these network traces. The queue length of Mahimahi is set to 60, and the drop-tail strategy is adopted.
@@ 5.2 Trade off between bitrate and quality
Prompt rank. We illustrate the variation in visual quality of Promptus under different ranks, as depicted in Figure 10. It indicates that the higher the prompt rank, the higher the quality of Promptus. For example, when the prompt rank increases from 4 to 16, the LPIPS decreases from 0.265 to 0.221 (on the line with a keyframe interval of 1). This is because the larger the rank, the stronger the representational capability of the prompt, allowing it to fit the target image more accurately, as described in §3.2. We present a visualization of the fitting results at different ranks in Figure 7. It indicates that as the rank increases, the prompt can fit more details. For instance, when the rank is 4, the earrings are lost. When the rank increases to 16, the earrings are successfully fitted.
Keyframe interval. Figure 10 also shows the impact of different keyframe intervals on quality. The smaller the keyframe interval, the higher the quality of Promptus. For example, when the prompt rank is 16, reducing the keyframe interval from 10 to 1 decreases the LPIPS from 0.274 to 0.221. This is because, when the keyframe interval is smaller, the distance between keyframes in the prompt space is smaller. At this time, the linear interpolation of keyframe prompts can more accurately approximate the prompts of intermediate frames, as described in §3.3.
Figure 8: The impact of prompt rank and keyframe interval on visual quality. It shows that visual quality improves with increasing rank and decreasing keyframe interval.
Figure 9: Tradeoff between bitrate and quality for Promptus. The quality monotonically increases with the increase in bitrate. Besides, Promptus can achieve scalable bitrates.
Figure 10: Average frame quality on 4 datasets. It proves the generality of Promptus across domains. The more high-fidelity details a video has, the greater the advantage of Promptus.
@@
Figure 11: Frame quality CDF and mean of Promptus and baselines at four target bitrates. It indicates Promptus achieves better transmission efficiency across all bitrate levels. The lower the bitrate, the greater the advantage of Promptus. Detail in  §5.3.
Quality-bitrate tradeoff. Increasing the prompt rank and reducing the keyframe interval lead to a rapid rise in bitrate while improving quality. Therefore, we present the tradeoff between bitrate and quality, as shown in Figure 10. First, it illustrates that overall, the quality monotonically increases with the increase in bitrate. Thus, to optimize quality, Promptus sends prompts at a bitrate closest to the available bandwidth. Second, Promptus can achieve scalable bitrates. For example, by adjusting the rank in the range of 4 to 32 and the keyframe interval in the range of 2 to 8, Promptus’s bitrate spans from 113 kbps to 4284 kbps. Third, at the same bitrate, the quality varies for different configurations. For instance, when the bitrate is 550 kbps, the parameter configuration with a rank of 8 and a keyframe interval of 4 has an LPIPS of 0.255, while the parameter configuration with a rank of 16 and a keyframe interval of 8 has an LPIPS of 0.264, which is lower in quality than the former. So at the same bitrate, we tend to choose parameter configurations with smaller keyframe intervals.
5.3 Transmission efficiency
This section demonstrates the transmission efficiency. Figure 11 shows the CDF of the frame quality under 4 bitrate levels. Since a lower LPIPS represents better visual quality, a leftward shift of the curve in the figure represents more high-quality frames, and thus higher transmission efficiency. We also calculate the average LPIPS for each method, represented by the vertical lines in the figure.
@@ 5.3 Transmission efficiency
First, Promptus achieves better transmission efficiency across all bitrate levels. For example, in Figure 11, the curves of Promptus are all to the left of the baseline curves. Second, Promptus can achieve more than a 4x bandwidth reduction while preserving the same perceptual quality. For example, the average LPIPS of Promptus under 140 kbps is better than H.265 under 540 kbps. Third, the lower the bitrate, the greater the advantage of Promptus. At a high level, as the bitrate decreases from 540 kbps to 140 kbps, the distance between Promptus’s curves and the baselines’ curves gradually widens. At a low level, when the bitrate is 540 kbps, the average LPIPS of Promptus is 0.018, 0.085 and 0.037 lower than VAE, H.265 and VQGAN, respectively. When the bitrate is 140 kbps, this difference further increases to 0.139, 0.118 and 0.046. This is because when the bitrate is reduced, the traditional codecs use coarser quantization and lose many high-frequency details, resulting in blurriness and block artifacts in the video, which significantly impairs the perceptual quality. VAE mitigates this phenomenon by mapping images to a low-dimensional latent space, reserving more bitrate for video encoding. Therefore, at most bitrates, the quality of VAE is superior to H.265. However, at extremely low bitrates (such as 140 kbps), the latent space inevitably introduces distortions caused by video coding. At this point, the VAE Decoder introduces a large number of errors when reconstructing the images, causing a significant decrease in VAE’s quality. On the other hand, when the bitrate is reduced, Promptus reduces the representational capability of the prompt rather than degrading the video quality. This prevents Promptus from accurately describing the video content, resulting in slight misalignments in the generated frames. However, thanks to the inherent image generation capability of Stable Diffusion, Promptus’s frames still have good sharpness and details, thus outperforming the baselines. When the bitrate decreases, VQGAN uses fewer tokens, which also reduces its representational capability. However, it does not have gradient descent for quality compensation, making it inferior to Promptus.
5.4 Generality on different domains
@@
This section proves the generality of Promptus across different domains. Figure 10 shows the average frame quality for Promptus and four baselines on four datasets.
@@
First, Promptus has good transmission efficiency on different domains. As shown in Figure 10, Promptus achieves lower average LPIPS compared to most baselines on each dataset. To intuitively demonstrate this gain, we visualize the compression results of each method on the four datasets at low bitrates (such as 225 kbps) in Figure 12. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
Figure 12: Visualization of the compression results on different datasets. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details.
@@
Second, the more high-frequency details a video has, the greater the advantage of Promptus. For example, for the Animerun dataset with fewer details, the LPIPS of Promptus is 0.015 lower than H.265, which is not a significant advantage. However, for the detail-rich UVG, this difference further expands to 0.121. This is because for 2D animations with large areas of solid colors and simple details, H.265’s intra-frame prediction, block partitioning, and motion compensation techniques can handle them well, so Promptus’s performance gain is small. This is also why Promptus performs worse than NAS on this dataset, as simple 2D animations are the domain where super-resolution performs well. For detail-rich real-world videos, H.265 discards more high-frequency information during compression, thus damaging the perceptual quality. Although Promptus also loses high-frequency information, Stable Diffusion completes some lost information based on prior knowledge during generation, resulting in a smaller quality loss.
Figure 13: Fitting results for challenging examples, with a bitrate of 225 kbps. Although the Stable Diffusion itself struggles to generate specific elements, such as fingers and text, Promptus is able to fit them well. This is because end-to-end gradient descent fitting can compensate for the limitations of the Stable Diffusion itself.
Since Stable Diffusion itself struggles to generate specific elements, such as fingers and text. To further demonstrate the generality of Promptus, we also show its fitting results for these challenging examples in Figure 13. The results indicate that Promptus is able to fit them well. This is because end-to-end gradient descent fitting can compensate for the limitations of the Stable Diffusion itself.
5.5 Temporal Consistency
This section evaluates the temporal consistency of the generated videos. As shown in Figure 14, we visualize the entire videos as space-time X-t slices (li2024generative, ), with three scanning lines at different heights marked on the left. Both of our method and H.265 are at a bitrate of 225 kbps.
Results demonstrate that although Prompus generates video frame by frame, it maintains good temporal consistency. First, it can be seen that our videos are temporally aligned with the ground truth videos in terms of motion. Second, our videos exhibit almost no flickering phenomena. For example, pixels in the X-t slices mostly move smoothly along the t dimension. This is because the proposed interpolation-aware fitting (§3.3) allows each frame to be as similar as possible to the ground truth, and video deflickering lei2023blind is also used to further alleviate inconsistency between frames. Third, compared to H.265, our method reduces visual degradation caused by blocking artifacts. As shown in the red box of Figure 14, H.265 exhibits some zig-zag phenomena in X-t slices, while our method is free from these artifacts. We provide some demo videos in our open-source repository to intuitively showcase the quality.
Figure 14: To evaluate the temporal consistency of the generated videos, we show the videos second by second. We also visualize the entire videos as space-time X-t slices (li2024generative, ), with three scanning lines at different heights marked on the left. The results indicate that our videos are temporally aligned with the ground truth videos in terms of motion, and exhibit almost no flickering phenomena. More details are in §5.5.
5.6 Flexible Resolution
Figure 15: Promptus supports flexible resolutions (§5.6).
This section shows that Promptus supports flexible resolutions. As illustrated in Figure 15, our generated images exhibit near-perfect visual alignment with ground truth across all resolutions. In fact, the resolutions supported by Promptus depend on the generator employed. Since Stable Diffusion supports flexible resolutions, Promptus naturally inherits this characteristic. This also explains a potential limitation: Promptus performs poorly on low-resolution videos because Stable Diffusion cannot generate low-resolution images sdturbo . For instance, as shown in Figure 15, videos with 256*256 resolution achieved the lowest PSNR.